OpenStudy (idealist10):

Solve y'=(y^2)(e^-x)+4y+2e^x explicitly.

OpenStudy (idealist10):

@amistre64 @ganeshie8 @iambatman @myininaya

ganeshie8 (ganeshie8):

lookup bernouli equations

OpenStudy (idealist10):

I know the bernouli method but I don't know what v should be.

ganeshie8 (ganeshie8):

wait, this is not exactly bernouli y'-4y = (y^2)(e^-x)+2e^x

ganeshie8 (ganeshie8):

we can't factor out y^2 on right hand side so this is not bernouli yet

OpenStudy (idealist10):

So what should we do?

OpenStudy (idealist10):

@myininaya @iambatman @ash2326

OpenStudy (anonymous):

Hint: it's a quadratic in disguise.

OpenStudy (anonymous):

The right side, I mean.

OpenStudy (anonymous):

Well maybe it's not exactly a hint, just an observation :P

OpenStudy (anonymous):

There are some initial steps you can take to make this a semi-homogeneous equation. Let \(u=e^x\). From the original eq, factor out \(e^x\): \[\begin{align*} y'&=e^x\left(y^2e^{-2x}+4ye^{-x}+2\right)\\\\ y'&=u\left(\left(\frac{y}{u}\right)^2+4\frac{y}{u}+2\right) \end{align*}\] Let \(y=ut\), so \(y'=ut'+t\): \[\begin{align*} ut'+t&=u\left(t^2+4t+2\right)\\\\ t'&=t^2+\left(4-\frac{1}{u}\right)t+2 \end{align*}\] Hmm...

OpenStudy (idealist10):

Then what would you do?

OpenStudy (anonymous):

I should have stuck with my first intuition... @Idealist10 you have everything you need to solve with ganeshie's link.

OpenStudy (idealist10):

Thank you!